OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 8 and 15 units in length.

OpenStudy (ghazi):

square root of 15^2+ 8^2

OpenStudy (ghazi):

17

OpenStudy (ajprincess):

Use pythagoras's theorem \(a^2+b^2=c^2\) Here a=8 and b=15